In Desmos, graph y= (-5x, 4y)


  1. state the domain and range of the original function
  2. state the domain and range of the inverse function
  3. the inverse reflects over?
1
Expert's answer
2021-03-24T15:00:36-0400

Given,

"y= (-5x, 4y)"

We can write the given function as

"y=-5x" and "x= 4y"



in the above we can see that the dark red line is representing "y=-5x" and dark blue line is representing "4y=x"

Both the function is a linear function so the domain and range of both the function be,

"D=(-\infty,\infty)"

"R=(-\infty,\infty)"

c)

"y=-5x"

Now, taking the inverse,

"x=\frac{-y}{5}"

"f^{-1}(x)=\frac{-x}{5}"

Similarly for the second function,

"y=\frac{x}{4}"

taking the inverse,

"f^{-1}(x)=4x"

As both the function are linear, and a linear function will be invertible as long as it's non-constant.